# LIME: Link-based User-Item Interaction Modeling with Decoupled XOR Attention for Efficient Test Time Scaling

### v2

@@ 1 Introduction
Modern recommendation systems operate at a massive scale, facing the challenge of ranking millions of candidate items within strict real-time latency constraints. To succeed, ranking models must navigate a fundamental trade-off between computational efficiency and predictive accuracy. This has led to two dominant but conflicting architectural paradigms. On one end of the spectrum is the two-tower model (Covington et al., 2016a), which achieves unparalleled inference speed by encoding users and items into separate, independent representations. This separation enables the use of efficient approximate nearest-neighbor search (Shrivastava & Li, 2014), making real-time recommendations feasible at scale.
On the other end are powerful cross-attention Transformer models like SASRec (Kang & McAuley, 2018b), which deliver state-of-the-art accuracy by encoding users’ long interaction history (UIH) via multi-layer self attention and explicitly modeling deep, contextual interactions between that and each candidate item. Although this approach provides rich expressiveness, it comes with a significant computational cost: the self-attention over long sequences and the cross-attention between the full user sequence and each candidate become major performance bottlenecks. This architectural dilemma is becoming increasingly acute. The push for higher quality recommendations now demands scaling along three axes simultaneously: accommodating vast candidate sets, processing longer user histories, and deploying models of increasing complexity. These demands are fundamentally at odds, as existing efficient models lack expressiveness, while expressive models lack efficiency. Hybrid approaches (Li et al., 2022) offer only incremental improvements.
@@ 2.1 Efficient Ranking Architectures
The design of ranking models is dominated by a conflict between efficiency and interaction depth. On one end of the spectrum, two-tower models (Covington et al., 2016b; Yi et al., 2019) achieve unparalleled efficiency. By encoding users and items into separate embedding spaces, they enable fast candidate retrieval using Approximate Nearest Neighbor (ANN) search. However, this separation prevents deep, feature-level interactions, limiting model expressiveness. On the other end, cross-attention models like DIN (Zhou et al., 2018) enable rich, target-aware interactions by dynamically attending to user history for each candidate, but their per-item computational cost makes them prohibitive for ranking large candidate sets.
Several approaches have sought to bridge this gap. Hybrid models add shallow interaction layers on top of a two-tower base (Li et al., 2022), though these often provide only marginal improvements. LIME offers a more fundamental solution. It introduces a novel bridge built upon a fixed set of global, learnable parameters, which we term link embeddings. These link embeddings act as an intermediary, allowing LIME to capture the rich dynamics of cross-attention while retaining the architectural efficiency of a two-tower model. This design, where a static, user-independent key space (raw links) retrieves information from a dynamic, personalized value space (personalized links), is conceptually similar to key-value memory networks (Miller et al., 2016) and the retrieval stage of retrieval-augmented models, enabling massive pre-computation. This approach directly tackles the sequential nature of user history and the need for scalable target-item interaction, striking a new balance between efficiency and expressiveness.
@@ 2.2 Efficient Attention for Long Sequence Modeling
The Transformer’s quadratic complexity ($O(N^{2})$) for self-attention remains a fundamental bottleneck for modeling long sequences. This challenge has driven the recent advances in Large Language Models (LLMs), moving beyond early approximations such as kernelization (Choromanski et al., 2021) or low-rank projections (Wang et al., 2020). More recent breakthroughs include new model classes—State Space Models (e.g., Mamba (Gu & Dao, 2023))—and hardware-aware techniques such as Lightning Attention (Dao et al., 2024) that directly optimize attention computation.
Within recommender systems, progress has often focused on adapting NLP-inspired attention mechanisms to click-through rate (CTR) prediction (Zhang et al., 2022; Li et al., 2023; Song et al., 2025) or using pruning to shorten sequences before attention is applied (Pi et al., 2020). While general-purpose mechanisms for handling set-based inputs exist, they often do not align with the specific needs of sequential recommendation. For instance, the Set Transformer (Lee et al., 2019) and its Pooling Multi-Head Attention (PMA) mechanism efficiently summarize a set into a fixed-size representation using learnable seed vectors. However, their primary goal is permutation-invariant summarization, whereas ranking requires target-aware representations that preserve the sequential nature of user histories.
@@ 3.2 From LIME-MHA to LIME-XOR: Scaling to Deeper Interactions
where $M_{\text{xor}}[i,j]:=1_{i\in[0,|E|)}\wedge 1_{j\in[0,|E|)}$ is the exclusive-or mask pattern that ensures the source and target embeddings attend to one another only. As depicted in Figure 3, the XOR mask structurally eliminates the expensive history-to-history ($E\leftrightarrow E$) interactions. Instead, it facilitates an efficient, two-way, block-wise attention (7): the link embeddings attend to the user history ($L^{C}\to E$), and crucially, the user history (source) embeddings now also attend back to the link embeddings ($E\to L^{C}$).
@@ 3.3 LIME’s Architectural Advantages for Efficient Inference
To facilitate this efficiency, we employ a technique called decoupled attention, used in the final candidate interaction stage (Equation 4). Instead of using the user-specific personalized links ($L^{P}$) for both keys and values, we use the raw, user-independent link embeddings ($L$) as keys and the personalized links ($L^{P}$) as values. This asymmetric structure ensures that the expensive Query-Key dot product, $\phi(CL^{t})$, depends only on item-side information (candidate embeddings $C$ and raw links $L$). Consequently, this attention weight matrix can be pre-computed for all items in the corpus and stored in an efficient key-value store or index, such as FAISS (Johnson et al., 2017), effectively creating a cache of attention weights.
@@
Stage 1: Offline Item-Side Pre-computation.
This stage is performed offline whenever the model or item catalog is updated, eliminating redundant computation during online serving.
  • Item Embedding Generation: Standard item features are processed by an item tower to produce item embeddings.
  • QK Cache Generation: The decoupled attention weights ($\phi(CL^{\top})$) between all candidate item embeddings ($C$) and the raw link embeddings ($L$) are pre-computed and cached. This transforms the most intensive part of cross-attention into a simple lookup.
Stage 2: Online User-Side Computation.
This stage runs once per user request and is independent of the number of candidates being scored.
  • Personalized Link Embedding Generation: The user’s context and interaction history are processed by the LIME user module (using either MHA or the multi-layer XOR Transformer) to produce the personalized link embeddings ($L^{P}$), as described in Equations (2)–(8).
Crucially, this user-side computation can be executed in parallel with the candidate retrieval process. In a production environment, its latency is therefore largely masked, making even a deep, multi-layer Transformer on the user side feasible.
Stage 3: Lightweight User-Item Interaction.
This final stage is executed online for each candidate but is extremely lightweight.
  • Final Score Computation: For each candidate, the pre-computed attention weights are retrieved from the QK Cache. These weights are used to perform a simple weighted sum over the personalized link embeddings ($L^{P}$) to generate the final LIME embedding. This embedding is then passed to a shallow interaction network for scoring.
@@ 4 Experimental Results
4.1 Baselines
We test multiple baseline models ranging from the simplest sum-pooling of user history embeddings to other sequence compression techniques listed below.
TTSN (Two-tower sparse network): Depicted in Figure 2, this baseline applies sum-pooling of all the user history embeddings instead of target attention with candidate embeddings.
4.2 Skylines
We cannot launch naive target attention (candidate against user history) directly for long user histories and large numbers of candidates to rank. But such powerful models can serve as our skyline goal on performance.
MHA Skyline computes a single layer of target attention between candidate items and user history items.
HSTU Skyline computes 3 layers of causal self attention and cross attention at each layer, between the candidate items and user history items.
4.3 LIME Variants
To thoroughly evaluate our proposed architecture, we conducted experiments with two primary variants of LIME that differ specifically in the sophistication of the link personalization module.
LIME-MHA: This is the fundamental implementation of our model, as described in Equation (3). It employs a standard attention mechanism to generate a personalized representation for each link by pooling information from the user’s historical item embeddings.
LIME-XOR: This advanced variant, detailed in Equation (8), incorporates 3-layers of HSTU between the user history and link embeddings for deeper contextualization with our proposed efficient XOR-style masking.
@@ 4.4 Accuracy
Normalized Entropy (NE) (He et al., 2014) is defined for binary classification task as
$\rm{NE}(\{(p_{i},\ell_{i})\}_{i=1}^{n}):=\frac{\sum_{i=1}^{n}\ell_{i}\log p_{i}+(1-\ell_{i})\log(1-p_{i})}{\log(\sum_{i=1}^{n}\ell_{i}/n)}.$
Note that similar to logloss or binary cross entropy, lower NE means better accuracy. Usually $0.1\%$ drop in NE will lead to online metric improvement.
@@ 4.6 Public Experiments Setup
We also benchmark on public datasets, namely Taobao Ads (Lyu et al., 2020) and KuaiRand-1K dataset (Gao et al., 2022). Taobao-Ads contains 25 million interactions with a maximum sequence length of 50 whereas KuaiRand-1K contains 12 million interactions with a maximum sequence length of 256. To unify data processing and evaluation framework, we leaned heavily on FuxiCTR, a comprehensive sequential recommendation model benchmark platform (Zhu et al., 2021).
@@
We also benchmark existing sequence-compression methods, including Truncated MHA—which applies the MHA skyline only to the most recent $\ell$ interacted items—and Linformer-style LREA (Song et al., 2025) with a low rank of $\ell$. At the same low rank, LIME outperforms both compression techniques. Additionally we benchmark skyline models SASRec (Kang & McAuley, 2018a) (a state-of-the-art transformer) and DIN (Deep Interest Network) (Zhou et al., 2018) to validate our LIME variants and skylines are indeed improvements over existing state-of-the-art models.
@@ 4.7 Inference Speed
LIME is highly scalable for ranking large candidate sets, an advantage for pre-ranking and retrieval. Figure 1 demonstrates that while MHA and HSTU latency grows significantly with more candidates or longer user histories, LIME’s latency remains nearly constant (see Appendix F for a detailed comparison). This robustness makes it suitable for settings with very long sequences (e.g.,>>30k items). Furthermore, the user backbone computation can be parallelized with candidate retrieval, masking most of its latency in production and yielding even greater savings than shown in these benchmarks.
@@ C.1 Dataset Preparation
For the Taobao Ad dataset, we leverage the preprocessed version provided by the FuxiCTR (Zhu et al., 2021) sequential modeling platform. For the KuaiRand dataset (Gao et al., 2022), we preprocess it ourselves by partitioning the first 14 days of interactions for training and last 2 days for testing. We also discarded all items from the train and test set with less than 30 total interactions for more reliable results. For the industrial dataset, we took 3 days of logged data for model training and 6 hours of data for evaluation.
@@ Appendix G Scaling Law for LIME-XOR Decoupled Transformer
Appendix H LLM Usage in Preparation

### v3

@@ 1 Introduction
Modern recommendation systems operate at a massive scale, facing the challenge of ranking millions of candidate items within strict real-time latency constraints. To succeed, ranking models must navigate a fundamental trade-off between computational efficiency and predictive accuracy. This has led to two dominant but conflicting architectural paradigms. On one end of the spectrum is the two-tower model (Covington et al., 2016), which achieves unparalleled inference speed by encoding users and items into separate, independent representations. This separation enables the use of efficient approximate nearest-neighbor search (Shrivastava and Li, 2014), making real-time recommendations feasible at scale.
On the other end are powerful cross-attention Transformer models like SASRec (Kang and McAuley, 2018), which deliver state-of-the-art accuracy by encoding users’ long interaction history (UIH) via multi-layer self attention and explicitly modeling deep, contextual interactions between that and each candidate item. Although this approach provides rich expressiveness, it comes with a significant computational cost: the self-attention over long sequences and the cross-attention between the full user sequence and each candidate become major performance bottlenecks. This architectural dilemma is becoming increasingly acute. The push for higher quality recommendations now demands scaling along three axes simultaneously: accommodating vast candidate sets, processing longer user histories, and deploying models of increasing complexity. These demands are fundamentally at odds, as existing efficient models lack expressiveness, while expressive models lack efficiency. Hybrid approaches (Li et al., 2022) offer only incremental improvements.
@@ 2.1 Efficient Ranking Architectures
The design of ranking models is dominated by a conflict between efficiency and interaction depth. On one end of the spectrum, two-tower models (Covington et al., 2016; Yi et al., 2019) achieve unparalleled efficiency. By encoding users and items into separate embedding spaces, they enable fast candidate retrieval using Approximate Nearest Neighbor (ANN) search. However, this separation prevents deep, feature-level interactions, limiting model expressiveness. On the other end, cross-attention models like DIN (Zhou et al., 2018) enable rich, target-aware interactions by dynamically attending to user history for each candidate, but their per-item computational cost makes them prohibitive for ranking large candidate sets.
Several approaches have sought to bridge this gap. Hybrid models add shallow interaction layers on top of a two-tower base (Li et al., 2022), though these often provide only marginal improvements. LIME offers a more fundamental solution. It introduces a novel bridge built upon a fixed set of global, learnable parameters, which we term link embeddings. These link embeddings act as an intermediary, allowing LIME to capture the rich dynamics of cross-attention while retaining the architectural efficiency of a two-tower model. This design, where a static, user-independent key space (raw links) retrieves information from a dynamic, personalized value space (personalized links), is conceptually similar to key-value memory networks (Miller, 2016) and the retrieval stage of retrieval-augmented models, enabling massive pre-computation. This approach directly tackles the sequential nature of user history and the need for scalable target-item interaction, striking a new balance between efficiency and expressiveness.
@@ 2.2 Efficient Attention for Long Sequence Modeling
The Transformer’s quadratic complexity ($O(N^{2})$) for self-attention remains a fundamental bottleneck for modeling long sequences. This challenge has driven the recent advances in Large Language Models (LLMs), moving beyond early approximations such as kernelization (Choromanski et al., 2020) or low-rank projections (Wang et al., 2020). More recent breakthroughs include new model classes—State Space Models (e.g., Mamba (Gu and Dao, 2024))—and hardware-aware techniques such as Lightning Attention (Qin et al., 2024) that directly optimize attention computation.
Within recommender systems, progress has often focused on adapting NLP-inspired attention mechanisms to click-through rate (CTR) prediction (Qi et al., 2022; Chen et al., 2022; Song et al., 2025) or using pruning to shorten sequences before attention is applied (Pi et al., 2020). While general-purpose mechanisms for handling set-based inputs exist, they often do not align with the specific needs of sequential recommendation. For instance, the Set Transformer (Lee et al., 2019) and its Pooling Multi-Head Attention (PMA) mechanism efficiently summarize a set into a fixed-size representation using learnable seed vectors. However, their primary goal is permutation-invariant summarization, whereas ranking requires target-aware representations that preserve the sequential nature of user histories.
@@ 3.2 From LIME-MHA to LIME-XOR: Scaling to Deeper Interactions
where $M_{\text{xor}}[i,j]:=1_{i\in[0,|E|)}\wedge 1_{j\in[0,|E|)}$ is the exclusive-or mask pattern that ensures the source and target embeddings attend to one another only. As depicted in Figure 3, the XOR mask structurally eliminates the expensive history-to-history ($E\leftrightarrow E$) interactions. Instead, it facilitates an efficient, two-way, block-wise attention (7): the link embeddings attend to the user history ($L^{C}\to E$), and crucially, the user history (source) embeddings also attend back to the link embeddings ($E\to L^{C}$).
@@ 3.3 LIME’s Architectural Advantages for Efficient Inference
To facilitate this efficiency, we employ decoupled attention, used in the final candidate interaction stage (Equation 4). Instead of using the user-specific personalized links ($L^{P}$) for both keys and values, we use the raw, user-independent link embeddings ($L$) as keys and the personalized links ($L^{P}$) as values. This asymmetric structure ensures that the expensive Query-Key dot product, $\phi(CL^{t})$, depends only on item-side information (candidate embeddings $C$ and raw links $L$). Consequently, this attention weight matrix can be pre-computed for all items in the corpus and stored in an efficient key-value store or index, such as FAISS (Johnson et al., 2017), effectively creating a cache of attention weights.
@@
Offline Item-Side Pre-computation. This stage is performed offline whenever the model or item catalog is updated, eliminating redundant computation during online serving. Standard item features are processed by an item tower to produce item embeddings. The decoupled attention weights ($\phi(CL^{\top})$) between all candidate item embeddings ($C$) and the raw link embeddings ($L$) are pre-computed and cached. This transforms the most intensive part of cross-attention into a simple lookup.
Online User-Side Computation. This stage runs once per user request and is independent of the number of candidates being scored. The user’s context and interaction history are processed by the LIME user module (using either MHA or the multi-layer XOR Transformer) to produce the personalized link embeddings ($L^{P}$), as described in Equations (2)–(8). Crucially, this user-side computation can be executed in parallel with the candidate retrieval process. In a production environment, its latency is therefore largely masked, making even a deep, multi-layer Transformer on the user side feasible.
Lightweight User-Item Interaction. This final stage is executed online for each candidate but is extremely lightweight. For each candidate, the pre-computed attention weights are retrieved from the QK Cache. These weights are used to perform a simple weighted sum over the personalized link embeddings ($L^{P}$) to generate the final LIME embedding. This embedding is then passed to a shallow interaction network for scoring.
@@ 4 Experimental Results
4.1 Baselines, Skylines, and LIME Variants
For baseline we mainly considered TTSN (Two-tower sparse network): Depicted in Figure 2, this baseline applies sum-pooling of all the user history embeddings instead of target attention with candidate embeddings.
While we cannot launch naive target attention (candidate against user history) directly for long user histories and large numbers of candidates to rank. But such powerful models can serve as our skyline goal on performance. MHA Skyline computes a single layer of target attention between candidate items and user history items. HSTU Skyline computes 3 layers of causal self attention and cross attention at each layer, between the candidate items and user history items.
Finally, to thoroughly evaluate our proposed architecture, we conducted experiments with two primary variants of LIME that differ specifically in the sophistication of the link personalization module. LIME-MHA: This is the fundamental implementation of our model, as described in Equation (3). It employs a standard attention mechanism to generate a personalized representation for each link by pooling information from the user’s historical item embeddings. LIME-XOR: This advanced variant, detailed in Equation (8), incorporates 3-layers of HSTU between the user history and link embeddings for deeper contextualization with our proposed efficient XOR-style masking.
@@ 4.2 Accuracy
Normalized Entropy (NE) (He et al., 2014) is defined for binary classification task as $\rm{NE}(\{(p_{i},\ell_{i})\}_{i=1}^{n}):=\frac{\sum_{i=1}^{n}\ell_{i}\log p_{i}+(1-\ell_{i})\log(1-p_{i})}{\log(\sum_{i=1}^{n}\ell_{i}/n)}$. Note that similar to logloss or binary cross entropy, lower NE means better accuracy. Usually $0.1\%$ drop in NE will lead to online metric improvement.
@@ 4.4 Public Experiments Setup
We also benchmark on public datasets, namely Taobao Ads (Lyu et al., 2020) and KuaiRand-1K dataset (Gao et al., 2022). Taobao-Ads contains 25 million interactions with a maximum sequence length of 50 whereas KuaiRand-1K contains 12 million interactions with a maximum sequence length of 256. To unify data processing and evaluation framework, we leaned heavily on FuxiCTR, a comprehensive sequential recommendation model benchmark platform (Zhu et al., 2020).
@@
We also benchmark existing sequence-compression methods, including Truncated MHA—which applies the MHA skyline only to the most recent $\ell$ interacted items—and Linformer-style LREA (Song et al., 2025) with a low rank of $\ell$. At the same low rank, LIME outperforms both compression techniques. Additionally we benchmark skyline models SASRec (Kang and McAuley, 2018) (a state-of-the-art transformer) and DIN (Deep Interest Network) (Zhou et al., 2018) to validate our LIME variants and skylines are indeed improvements over existing state-of-the-art models.
@@
4.5 Ablation Study
4.6 Comparison to Linear Attention Methods
@@ 4.7 Inference Speed
Figure 1 demonstrates that while MHA and HSTU latency grows significantly with more candidates or longer user histories, LIME’s latency remains nearly constant (see Appendix F for a detailed comparison). This robustness makes it suitable for settings with very long sequences (e.g.,>>30k items). Furthermore, the user backbone computation can be parallelized with candidate retrieval, masking most of its latency in production and yielding even greater savings.
@@ 5 Analysis and Discussions
6 Limitations and Future Work
Despite LIME’s strong empirical performance, several limitations warrant discussion. First, LIME’s effectiveness relies on the low-rank approximation assumption; domains with highly fragmented user interests may require higher-rank representations. Second, pre-computing QK caches requires periodic updates for evolving item catalogs, which makes it better suited for early stage ranking.
We are actively exploring the following directions: (1) building a larger pool of link embeddings and dynamically selecting link subsets based on user-side features or context, enabling personalized link allocation across different user segments, (2) hierarchical link structures to capture interests at multiple granularities, (3) extensions to multi-modal recommendation and cross-domain transfer learning, and (4) interpretability analysis to understand what semantic patterns individual links capture.
@@ C.1 Dataset Preparation
For the Taobao Ad dataset, we leverage the preprocessed version provided by the FuxiCTR (Zhu et al., 2020) sequential modeling platform. For the KuaiRand dataset (Gao et al., 2022), we preprocess it ourselves by partitioning the first 14 days of interactions for training and last 2 days for testing. We also discarded all items from the train and test set with less than 30 total interactions for more reliable results. For the industrial dataset, we took 3 days of logged data for model training and 6 hours of data for evaluation.
@@ Appendix G Scaling Law for LIME-XOR Decoupled Transformer
Appendix H Derivations for XOR Attention Kernel
H.1 Forward Pass
Let the output of the XOR-attention be $O=(O[S],O[T])$, where $S,T$ stand for source and target respectively. In the context of LIME, source is the user history item embeddings, while target is the link embeddings. Similarly define $Q[S],Q[T],K[S],K[T],V[S],V[T]$ to be the source and target portion of the query, key, value embedding sequences.
$O[T]$ $=\phi(Q[T]K[S]^{\top})V[S]$
$O[S]$ $=\phi(Q[S]K[T]^{\top})V[T]$
H.2 Backward Pass
Let $dO[T]$ denote an infinitesimally small change in $O[T]$, also known as the its differential. Similarly define $dO[S],dV[S],dV[T],dQ[S],dQ[T],dK[S],dK[T]$.
Trivially we have
$\frac{\partial O[T]}{\partial Q[S]}$ $=\frac{\partial O[T]}{\partial K[T]}=\frac{\partial O[T]}{\partial V[T]}=0$
$\frac{\partial O[S]}{\partial Q[T]}$ $=\frac{\partial O[S]}{\partial K[S]}=\frac{\partial O[S]}{\partial V[S]}=0$
The total differential of the loss function is given by
$dL=\operatorname{\operatorname{Tr}}\left(\left(\frac{\partial L}{\partial O[S]}\right)^{\top}dO[S]\right)+\operatorname{\operatorname{Tr}}\left(\left(\frac{\partial L}{\partial O[T]}\right)^{\top}dO[T]\right).$
Since $dO[S]=\phi(Q[S]K[T]^{\top})dV[T]$ and $dO[T]=\phi(Q[T]K[S]^{\top})dV[S]$,
$dL_{V}=\operatorname{\operatorname{Tr}}\left(\left(\frac{\partial L}{\partial O[S]}\right)^{\top}\phi(Q[S]K[T]^{\top})dV[T]\right)+\operatorname{\operatorname{Tr}}\left(\left(\frac{\partial L}{\partial O[T]}\right)^{\top}\phi(Q[T]K[S]^{\top})dV[S]\right).$
By comparing coefficient with the matrix chain rule $dL_{V[T]}=\operatorname{\operatorname{Tr}}\left(\left(\frac{\partial L}{\partial V[T]}\right)^{\top}dV[T]\right)$, we see that
$\frac{\partial L}{\partial V[T]}=\phi(K[T]Q[S]^{\top})\frac{\partial L}{\partial O[S]}.$
Similarly
$\frac{\partial L}{\partial V[S]}=\phi(K[S]Q[T]^{\top})\frac{\partial L}{\partial O[T]}.$
Next for derivatives with respect to $Q$, we have
$dL_{Q[S]}$ $=\operatorname{\operatorname{Tr}}\left(\left(\frac{\partial L}{\partial O[S]}\right)^{\top}\phi^{\prime}(Q[S]K[T]^{\top})V[T]K[T]^{\top}dQ[S]\right)$
$dL_{Q[T]}$ $=\operatorname{\operatorname{Tr}}\left(\left(\frac{\partial L}{\partial O[T]}\right)^{\top}\phi^{\prime}(Q[T]K[S]^{\top})V[S]K[S]^{\top}dQ[T]\right).$
Hence
$\frac{\partial L}{\partial Q[S]}$ $=K[T]V[T]^{\top}\phi^{\prime}(K[T]Q[S]^{\top})\frac{\partial L}{\partial O[S]}$
$\frac{\partial L}{\partial Q[T]}$ $=K[S]V[S]^{\top}\phi^{\prime}(K[S]Q[T]^{\top})\frac{\partial L}{\partial O[T]}.$
Finally from
$dL_{K[T]}$ $=\operatorname{\operatorname{Tr}}\left(\left(\frac{\partial L}{\partial O[S]}\right)^{\top}\phi^{\prime}(Q[S]K[T]^{\top})V[T]Q[S]^{\top}dK[T]\right)$
$dL_{K[S]}$ $=\operatorname{\operatorname{Tr}}\left(\left(\frac{\partial L}{\partial O[T]}\right)^{\top}\phi^{\prime}(Q[T]K[S]^{\top})V[S]Q[T]^{\top}dK[S]\right),$
we get
$\frac{\partial L}{\partial K[S]}$ $=Q[T]V[S]^{\top}\phi^{\prime}(K[S]Q[T]^{\top})\frac{\partial L}{\partial O[T]}$
$\frac{\partial L}{\partial K[T]}$ $=Q[S]V[T]^{\top}\phi^{\prime}(K[T]Q[S]^{\top})\frac{\partial L}{\partial O[S]}$
Appendix I Triton PseudoCode
Algorithm 1 XOR Mask and Denominator Computation
1:Query index $i$, key index $j$, number of sources $n_{s}$
2:Binary mask $m$, normalization denominator $d$
[…]
Algorithm 2 Forward Pass with Block Range Selection
[…]
2:Output $\mathbf{O}\in\mathbb{R}^{n\times d}$
3:for $q_{\text{start}}\leftarrow 0$ to $n$ step $B_{M}$ do $\triangleright$ Parallel over query blocks
4:  $\mathbf{Q}_{b}\leftarrow\mathbf{Q}[q_{\text{start}}:q_{\text{start}}+B_{M},:]$ $\triangleright$ Load to SRAM
5:  $\text{acc}\leftarrow\mathbf{0}_{B_{M}\times d}$
6:
7:  // Block range selection (critical optimization)
8:  if $q_{\text{start}}+B_{M}\leq n_{s}$ then $\triangleright$ Pure source queries
9:   $k_{\text{lo}},k_{\text{hi}}\leftarrow n_{s},n$ $\triangleright$ Load target keys only
10:  else if $q_{\text{start}}\geq n_{s}$ then $\triangleright$ Pure target queries
11:   $k_{\text{lo}},k_{\text{hi}}\leftarrow 0,n_{s}$ $\triangleright$ Load source keys only
[…]
13:   $k_{\text{lo}},k_{\text{hi}}\leftarrow 0,n$
14:  end if
15:
16:  for $k_{\text{start}}\leftarrow k_{\text{lo}}$ to $k_{\text{hi}}$ step $B_{N}$ do $\triangleright$ Over selected K,V blocks
17:   $\mathbf{K}_{b}\leftarrow\mathbf{K}[:,k_{\text{start}}:k_{\text{start}}+B_{N}]$ $\triangleright$ Load K,V to SRAM
18:   $\mathbf{V}_{b}\leftarrow\mathbf{V}[k_{\text{start}}:k_{\text{start}}+B_{N},:]$
19:   $\mathbf{S}\leftarrow\mathbf{Q}_{b}\mathbf{K}_{b}$ $\triangleright$ Compute scores (tensor cores)
20:
21:   for $i\leftarrow 0$ to $B_{M}$ do $\triangleright$ Mask & normalize in registers
22:     for $j\leftarrow 0$ to $B_{N}$ do
23:      $m,d\leftarrow\textsc{XorMask}(q_{\text{start}}+i,k_{\text{start}}+j,n_{s})$
24:      $\mathbf{S}[i,j]\leftarrow m\ ?\ \text{SiLU}(\mathbf{S}[i,j])/d:0$
25:     end for
26:   end for
27:
28:   $\text{acc}\leftarrow\text{acc}+\mathbf{S}\mathbf{V}_{b}$ $\triangleright$ Accumulate (tensor cores)
29:  end for
30:
31:  $\mathbf{O}[q_{\text{start}}:q_{\text{start}}+B_{M},:]\leftarrow\text{acc}$ $\triangleright$ Write to HBM
32:end forreturn $\mathbf{O}$
Algorithm 3 Backward Pass (Transposed Access Pattern)
[…]
2:Gradients $\mathbf{dQ},\mathbf{dK},\mathbf{dV}$
[…]
4:for $k_{\text{start}}\leftarrow 0$ to $n$ step $B_{N}$ do $\triangleright$ Parallel over K,V blocks
5:  $\mathbf{K}_{b}\leftarrow\mathbf{K}[:,k_{\text{start}}:k_{\text{start}}+B_{N}]$ $\triangleright$ Load K,V to SRAM (resident)
6:  $\mathbf{V}_{b}\leftarrow\mathbf{V}[k_{\text{start}}:k_{\text{start}}+B_{N},:]$
7:  $\mathbf{dK}_{\text{acc}},\mathbf{dV}_{\text{acc}}\leftarrow\mathbf{0}_{d\times B_{N}},\mathbf{0}_{B_{N}\times d}$
8:
9:  // Symmetric block range selection
10:  if $k_{\text{start}}+B_{N}\leq n_{s}$ then $\triangleright$ Source K,V block
11:   $q_{\text{lo}},q_{\text{hi}}\leftarrow n_{s},n$ $\triangleright$ Process target queries only
12:  else if $k_{\text{start}}\geq n_{s}$ then $\triangleright$ Target K,V block
13:   $q_{\text{lo}},q_{\text{hi}}\leftarrow 0,n_{s}$ $\triangleright$ Process source queries only
[…]
15:   $q_{\text{lo}},q_{\text{hi}}\leftarrow 0,n$
16:  end if
17:
18:  for $q_{\text{start}}\leftarrow q_{\text{lo}}$ to $q_{\text{hi}}$ step $B_{M}$ do $\triangleright$ Over selected Q blocks
19:   $\mathbf{Q}_{b}\leftarrow\mathbf{Q}[q_{\text{start}}:q_{\text{start}}+B_{M},:]$
20:   $\mathbf{dO}_{b}\leftarrow\mathbf{dO}[q_{\text{start}}:q_{\text{start}}+B_{M},:]$
21:   $\mathbf{S}\leftarrow\mathbf{Q}_{b}\mathbf{K}_{b}$ $\triangleright$ Recompute forward (activation remat)
22:
23:   for $i\leftarrow 0$ to $B_{M}$, $j\leftarrow 0$ to $B_{N}$ do
24:     $m,d\leftarrow\textsc{XorMask}(q_{\text{start}}+i,k_{\text{start}}+j,n_{s})$
25:     if $m$ then
26:      $s_{ij}\leftarrow\text{SiLU}(\mathbf{S}[i,j])/d$
27:      $\mathbf{S}[i,j]\leftarrow s_{ij}\cdot(1-s_{ij})\cdot(1+\mathbf{S}[i,j])/d$ $\triangleright$ SiLU gradient
28:     else
29:      $\mathbf{S}[i,j]\leftarrow 0$
30:     end if
31:   end for
32:
33:   $\mathbf{dV}_{\text{acc}}\leftarrow\mathbf{dV}_{\text{acc}}+\mathbf{S}^{\top}\mathbf{dO}_{b}$ $\triangleright$ Accumulate (tensor cores)
34:   $\mathbf{dK}_{\text{acc}}\leftarrow\mathbf{dK}_{\text{acc}}+(\mathbf{S}^{\top}\mathbf{Q}_{b})^{\top}$
35:   AtomicAdd($\mathbf{dQ}[q_{\text{start}}:q_{\text{start}}+B_{M},:]$, $(\mathbf{dO}_{b}\mathbf{V}_{b}^{\top})\mathbf{S}^{\top}$)
36:  end for
37:
38:  $\mathbf{dK}[:,k_{\text{start}}:k_{\text{start}}+B_{N}]\leftarrow\mathbf{dK}_{\text{acc}}$ $\triangleright$ Write gradients
39:  $\mathbf{dV}[k_{\text{start}}:k_{\text{start}}+B_{N},:]\leftarrow\mathbf{dV}_{\text{acc}}$
40:end forreturn $\mathbf{dQ},\mathbf{dK},\mathbf{dV}$
Appendix J LLM Usage in Preparation